OpenStudy (anonymous):

I dont get how to do this pls help Classify each equation as a linear equation in standard form, a linear equation not in standard form, or a nonlinear equation. Some problems on the bottom 1)2x+5y=2 2)xy=9 3)x+y=x-y

OpenStudy (anonymous):

1. linear equation not in standard form(stand form is y =mx + b) 2. nonlinear 3. linear equation not in standard for

OpenStudy (anonymous):

Then can you do these 4.x^2+y^2=16 5.2/x+3/y=3 6.x/4+y/7=1

OpenStudy (anonymous):

4. nonlinear 5. nonlinear 6. linear but not in standard form